Subject: Thinfilms morning report 2/24/05
From: james haydon <jhaydon@snf.stanford.edu>
Date: Thu, 24 Feb 2005 11:07:47 -0800
Thu, 24 Feb 2005 11:07:47 -0800
Thinfilms Morning Report 2/24/2005

 

 

Gryphon:            ok

Innotec;            ion gauge controller not working I should have the 
parts tomorrow innotec now has a leak that I will not be able to check 
out until next week.

STS Dep:            ok

Implanter:            noisy signal still but ion gauge problem is fixed 
if you look at the end station ion gauge I will try to get to the noise 
problem next week

Metalica:            ok

SCT:                ok

CPD:                ok                   

WbGen-CTB:            ok

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            ok

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok

 



Thinfilms Morning Report 2/24/2005

 

 

Gryphon:            ok

Innotec;            ion gauge controller not working I should have the parts tomorrow innotec now has a leak that I will not be able to check out until next week.

STS Dep:            ok

Implanter:            noisy signal still but ion gauge problem is fixed if you look at the end station ion gauge I will try to get to the noise problem next week

Metalica:            ok

SCT:                ok

CPD:                ok                   

WbGen-CTB:            ok

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            ok

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok